A motor car starts with the speed of 70 km/hr with its speed increasing every two hours by 10 kmph. In how many hours will it cover 345 kms?


[A] 2 (1/4) hrs


[E] None of these


[C] 4(1/2) hrs


[B] 4 hrs 5 min


 


[D] Can not be determined 

Grade:9

3 Answers

jhansi adapaka
10 years ago

it take aproximately 4 hours 40 min

Murtaza Khanjiwala
10 years ago

V-U=at

10=a(2)

a=5

S=ut+at2/2

s=(70)t+(5)t2/2

69=(14)t+t^2/2

0=t2+28t-138

solving this eqn we get t=4.27hrs.

raj
5 years ago
increasing speed in 2 hrs= 10km/hr
‘’                  ‘’      ‘’    1 hr= 5 km/hr 
 
extended speed = 70+5=75 km/hr 
time required= (345/75) = 4.6 hr